# Diagonalization Procedure for Reducible Massless ${\cal N}=1$ Supermultiplets

### v2

@@ 1 Introduction
There are various approaches to supersymmetric higher spin theories both on flat and on anti-de Sitter ($AdS$) backgrounds (see [1][5] for a pioneering work in this direction, and [6][10] for reviews). In most cases one considers irreducible representations of the Poincaré or $AdS_{d}$ groups111 A possible implication of ${\cal N}=2$ supersymmetry in higher spin theories on $dS_{4}$ background was recently discussed in [11]., which in turn form irreducible, and in some cases infinite-dimensional, representations of the corresponding supersymmetry algebra. On the other hand, consideration of reducible representations might open some new interesting possibilities. Systems with reducible representations of the Poincaré group were studied in detail in [12][15] and called triplets and generalized triplets, in case fields with mixed symmetries are included. The main advantage of considering reducible representations is that the corresponding free Lagrangians and supersymmetry transformations have a relatively simple form when compared to their irreducible counterparts. This simplicity of the quadratic Lagrangians usually proves to be very helpful when constructing cubic, and higher order, interactions, both for massive and massless higher spin fields [16][19], [20].
@@
Following this line of thought, we elaborate on the ${\cal N}=1$ supersymmetric model in $d=4,6,10$, space-time dimensions, constructed in [21][22]. In this superstring-inspired model, the fermionic sector consists of the fields described by totally symmetric spin–tensors, whereas the bosonic sector consists of both totally symmetric fields and fields with mixed symmetry of the simplest type [^2] Supersymmetry in $d$ dimensions in the unfolded formulation was recently considered in [23]. $Y(n-1,1)$. Because of the gauge invariant description, the system contains, along with physical fields, also fields which are gauge degrees of freedom both in bosonic and in fermionic sectors. One can show that, after the complete fixing of the gauge, the spectrum contains only massless reducible representations of the Poincaré group described by the Young diagrams $Y(n,0)$ and $Y(n-1,1)$ in the bosonic sector and Young diagrams $Y(n-1,0)$ in the fermionic sector. Both bosonic and fermionic representations are reducible, since the traces of the fields are non-zero on-shell. Moreover, the corresponding ${\cal N}=1$ supermultiplets are irreducible in $d=10$, but reducible in $d=4$ and $d=6$ also from the point of view of representations of supersymmetry algebra. As mentioned above, although this approach contains more fields compared to an approach based on irreducible representations of the supersymmetry and Poincaré groups, it allows one to construct Lagrangians for both higher and lower spin fields in a relatively simple form. For example, in this way one can obtain a new description of linearized ${\cal N}=1$ supergravity and linear supermultiplet in four dimensions, linearized ${\cal N}=1$ supergravity and $(1,0)$ tensor multiplet in six dimensions, and $N=1$ linearized supergravity in ten dimensions (see also [24] for a formulation of linearized $d=6$ supergravity in terms of reducible fields).
@@
The purpose of the present paper is to develop a procedure for the decomposition of fermionic and bosonic Lagrangians constructed in [21] into a direct sum of Lagrangians for individual irreducible representations of the Poincaré group [25][26]. This procedure will henceforth be referred to as a diagonalization procedure. For totally symmetric representations of the Poincaré group the diagonalization has been done in [27][29]. Here we show how to perform the diagonalization procedure for fermionic totally symmetric fields and for the fields with the mixed symmetry [^3] See [30][36] for different quadratic actions for mixed symmetry fields on Minkowski background. of the type $Y(n,1)$. Let us note, that without requiring of ${\cal N}=1$ supersymmetry, both bosonic and fermionic Lagrangians as well as the corresponding diagonalization procedures are valid for any ($d\geq 4$) number of space-time dimensions.
@@
The motivation behind developing the diagonalization procedure is two-fold. Firstly, although it is more convenient to construct cubic and higher-order interaction vertices for reducible representations, in certain cases it might still be preferable to consider the interactions between reducible modes. The second reason is the possibility of addressing the question of a modification of the system constructed [21] to an $AdS_{d}$ background [^4] The problem of a formulation of higher spin theories on backgrounds which are different from flat and $(A)dS_{d}$ spaces has been recently addressed in [37].. A deformation of bosonic and fermionic triplets that contain only totally symmetric fields to $AdS_{d}$ [13], [38][39], as well as the relevant diagonalization procedure [28] are known and can be applied directly to the present analysis. However, the presence of mixed-symmetry fields in the spectrum causes some difficulties, and we shall comment on them at the end of Section 4.
@@
In Section 3 we demonstrate in detail the procedure of diagonalization for the simplest case of $n=1$, when the physical fields in the bosonic sector are the second-rank symmetric and antisymmetric tensors, whereas in the fermionic sector one has a first-rank spinor-tensor. We also discuss various ways of deformation of this system to an $AdS_{d}$ background.
In Section 4 we repeat this procedure in the more complicated case of $n=2$. We show that after the diagonalization of the bosonic Lagrangian one obtains the Fronsdal Lagrangian for the spin $3$ field, i.e., for the irreducible representation $Y(3,0)$, a Lagrangian for the mixed symmmetry irreducible representation of type $Y(2,1)$, and two Lagrangians for vector fields, which correspond to the traces of above mentioned (initially reducible) representations. In the fermionic sector one obtains the fields with spins $\frac{5}{2}$, $\frac{3}{2}$ and $\frac{1}{2}$. The detailed study of these examples suggests a clear pattern of how to generalize the procedure to the higher-rank tensors, which is discussed in Section 5.
@@ 2 Reducible ${\cal N}=1$ supermultiplets
Let us briefly summarize the main features of the ${\cal N}=1$ supersymmetric system constructed in [21]. To this end, it is convenient to introduce an auxiliary Fock space, spanned by one bosonic and one fermionic auxiliary oscillators
@@
Let us consider the bosonic fields first, were one has a physical field $|\phi\rangle$ and a set of auxiliary fields. The fields are of type $|X^{(p,q)}\rangle$, where $p$ is a number of the bosonic oscillators and $q$ which can be either $1$ or 00, is the number of fermionic oscillators. Obviously the fields are symmetrical with respect to $p$-indices and no symmetry between $p$-indices and $q$-index is implied.
@@
where $|M\rangle=|B\rangle-i|A\rangle$. It can be shown (see [21] for the details), that one can use the gauge invariance (2.6) and the equations of motion in order to eliminate on–shell all fields in (2.2)–(2.3), except $|\phi\rangle$. This vector obeys massless Klein-Gordon equation, is transverse with respect to both sets of its indices and contains the totally symmetric representations of the Poincaré group of type $Y(n+1,0)$, fields with mixed symmetry of type $Y(n,1)$, as well as their traces.
The states in the fermionic sector contain only the bosonic oscillators, defined in (2.1). The fields, as well as a Fock vacuum carry the corresponding $d$dimensional spinorial index, which will not be written explicitly. The field content is
@@
where $|\xi^{(n-1,0)}\rangle$ is a parameter. Again, one can show, that it is possible to gauge away the non-physical fields $|\chi\rangle$ and $|\Sigma\rangle$, and one is left on-shell with the physical field $|\Psi\rangle$. This field obeys massless Dirac equation, is transverse and describes a massless field with spin $n+\frac{1}{2}$ as well as lower spins $n-\frac{1}{2},...,\frac{1}{2}$, since no zero gamma–trace condition is implied.
@@
which leaves the sum of the Lagrangians (2) and (2) invariant. One can check that the supersymmetry algebra close on-shell in the dimensions $d=4,6$ and $10$. Expressing the auxiliary fields via their own equations of motion (2.8), one can can write the supersymmetry transformation rules for fermions as
@@
Let us note that, the closure of the supersymmetry algebra, as well as the consistency of the equations (2.8) with the transformation rules (2.13)– requires the use of fermionic field equations, derived from the Lagrangian (2), as it should have been expected since we are considering on-shell supersymmetry.
@@ 3.1 Flat space-time
For the fermionic fields, the first two steps of the diagonalization procedure are not needed and one starts from the third step. Similarly to the case of the bosonic fields we introduce a spin $\frac{3}{2}$ field $\tilde{\Psi}_{\mu}$ and a gauge invariant spin $\frac{1}{2}$ field $\tilde{\Psi}$ as
@@
where the coefficient in front of the last term in the definition of the field $\tilde{\Psi}_{\mu}$ is chosen in such a way, that the Lagrangian (3.9) written in terms of the new fields is a sum of Fang – Fronsdal Lagrangians for spins $\frac{3}{2}$ and $\frac{1}{2}$
@@
Finally, using the equations (3.6), (3.7), (3.11), as well as (2.14) and (2.15), it is straightforward to rewrite the supersymmetry transformations in terms of irreducible modes. From the field content given above, one can see that this system describes a linearized ${\cal N}=1$ supergravity and the linear multiplet in case of $d=4$, a linearized ${\cal N}=1$ supergravity and $(1,0)$ tensor multiplet in case of $d=6$ and a linearized ${\cal N}=1$ supegravity in case of $d=10$ dimensions [21].
@@ 3.2 Deformation to $AdS$ space
In order to perform a deformation of the bosonic (2) and fermionic (2) Lagrangians to an $AdS_{d}$ background, one can proceed in two different ways. The first option is to start with the already decomposed system described by a sum of the Lagrangians (3.5), (3.8) and (3.12). Then we perform its deformation to $AdS_{d}$ space by replacing the usual partial derivatives with the covariant ones and adding the corresponding mass-like terms in order to keep all gauge symmetries intact. The Lagrangians for totally symmetric fermionic and bosonic fields can be directly read off from the equations (A) and (A) which we give here for convenience
@@
where $H_{\mu_{1}\mu_{2}\mu_{3}}=\partial_{[\mu_{1}}{\cal B}_{\mu_{2}\mu_{3}]}$. In case of $d=4$, one has another option for $AdS$ deformation. Namely, one can first dualize the field ${\cal B}_{\mu\nu}$ to a scalar $a(x)$ as $\partial_{\mu}a\sim\epsilon_{\mu\nu\rho\sigma}H^{\nu\rho\sigma}$ and then perform the $AdS_{4}$ deformation to a conformal scalar, with the mass–like term being the same as in (3.2) for the scalar field $\Phi$. As a result, one obtains either the linear ${\cal N}=1$ supermultiplet [40][41], or the Wess–Zumino supermultiplet, where the spin $\frac{1}{2}$ field being $\tilde{\Psi}$. The fields ${\Phi}_{\mu\nu}$ and $\tilde{\Psi}_{\mu}$, in turn form a supermultiplet of the ${\cal N}=1$ supergravity on $AdS_{4}$, constructed in[42].
@@
Using the commutation relations given in the Appendix, one can check, that Lagrangian (3.2) is invariant under the gauge transformations
@@
$\delta C_{\nu}=-\Box\lambda_{\nu}+\frac{1}{L^{2}}(d-1)\rho_{\mu}+\nabla_{\nu}\tau,$ (3.17)

### v3

@@ 1 Introduction
There are various approaches to supersymmetric higher spin theories both on flat and on anti-de Sitter ($AdS$) backgrounds (see [1][5] for a pioneering work in this direction, and [6][10] for reviews). In most cases one considers irreducible representations of the Poincaré or $AdS_{d}$ groups111 A possible implication of ${\cal N}=2$ supersymmetry in higher spin theories on $dS_{4}$ background was recently discussed in [11]., which in turn form irreducible, and in some cases infinite-dimensional, representations of the corresponding supersymmetry algebra. On the other hand, consideration of reducible representations might open some new interesting possibilities. Systems with reducible representations of the Poincaré group were studied in detail in [12][15] and called either triplets or generalized triplets, in the case fields with mixed symmetries are included. The main advantage of considering reducible representations is that the corresponding free Lagrangians and supersymmetry transformations have a relatively simple form when compared to their irreducible counterparts. This simplicity of the quadratic Lagrangians usually proves to be very helpful when constructing cubic and higher order interactions, both for massive and massless higher spin fields [16][19], [20].
@@
Following this line of thought, we elaborate on the ${\cal N}=1$ supersymmetric model in $d=4,6,10$, space-time dimensions, constructed in [21][22]. In this superstring-inspired model, the fermionic sector consists of the fields described by totally symmetric spin–tensors, whereas the bosonic sector consists of both totally symmetric fields and fields with mixed symmetry of the simplest type [^2] Supersymmetry in $d$ dimensions in the unfolded formulation was recently considered in [23]. $Y(n-1,1)$. Because of the gauge invariant description, the system contains, along with physical fields, also fields that are gauge degrees of freedom both in bosonic and in fermionic sectors. One can show that, after the complete fixing of the gauge, the spectrum contains only massless reducible representations of the Poincaré group described by the Young diagrams $Y(n,0)$ and $Y(n-1,1)$ in the bosonic sector and Young diagrams $Y(n-1,0)$ in the fermionic sector. Both bosonic and fermionic representations are reducible, since the traces of the fields are non-zero on-shell. Moreover, the corresponding ${\cal N}=1$ supermultiplets are irreducible in $d=10$, but reducible in $d=4$ and $d=6$ also from the point of view of representations of supersymmetry algebra. As mentioned above, although this approach contains more fields compared to an approach based on irreducible representations of the supersymmetry and Poincaré groups, it allows one to construct Lagrangians for both higher and lower spin fields in a relatively simple form. For example, in this way one can obtain: a new description of linearized ${\cal N}=1$ supergravity and a linear supermultiplet in four dimensions; linearized ${\cal N}=1$ supergravity and a $(1,0)$ tensor multiplet in six dimensions and $N=1$ linearized supergravity in ten dimensions (see also [24] for a formulation of linearized $d=6$ supergravity in terms of reducible fields).
@@
The purpose of the present paper is to develop a procedure for the decomposition of fermionic and bosonic Lagrangians constructed in [21] into a direct sum of Lagrangians for individual irreducible representations of the Poincaré group [25][26]. This procedure will henceforth be referred to as a diagonalization procedure. For totally symmetric representations of the Poincaré group the diagonalization has been carried out in [27][29]. Here we show how to perform the diagonalization procedure for fermionic totally symmetric fields and for the fields with mixed symmetry [^3] See [30][36] for different quadratic actions for mixed symmetry fields on Minkowski background. of the type $Y(n,1)$. Let us note, that if ${\cal N}=1$ supersymmetry is not required, both bosonic and fermionic Lagrangians, as well as the corresponding diagonalization procedures, are valid for any ($d\geq 4$) number of space-time dimensions.
@@
The motivation behind developing the diagonalization procedure is two-fold. Firstly, although it is more convenient to construct cubic and higher-order interaction vertices for reducible representations, in certain cases it might still be preferable to consider the interactions between reducible modes. The second reason is the possibility of addressing the question of a modification of the system constructed [21] to an $AdS_{d}$ background [^4] The problem of a formulation of higher spin theories on backgrounds which are different from flat and $(A)dS_{d}$ spaces has been recently addressed in [37].. Both deformation of bosonic and fermionic triplets that contain only totally symmetric fields to $AdS_{d}$ [13], [38][39], as well as the relevant diagonalization procedure [28], are known and can be applied directly to the present analysis. However, the presence of mixed-symmetry fields in the spectrum causes some difficulties, and we shall comment on these at the end of Section 4.
@@
In Section 3 we demonstrate in detail the diagonalization procedure for the simplest case of $n=1$, where the physical fields in the bosonic sector are second-rank symmetric and antisymmetric tensors, whereas in the fermionic sector one has first-rank spinor-tensor. We also discuss various ways to deform this system to an $AdS_{d}$ background.
In Section 4 we repeat this procedure in the more complicated case of $n=2$. We show that after diagonalization of the bosonic Lagrangian one obtains: the Fronsdal Lagrangian for the spin $3$ field, i.e., for the irreducible representation $Y(3,0)$; a Lagrangian for the mixed symmetry irreducible representation of type $Y(2,1)$; and two Lagrangians for vector fields, which correspond to the traces of abovementioned (initially reducible) representations. In the fermionic sector one obtains the fields with spins $\frac{5}{2}$, $\frac{3}{2}$ and $\frac{1}{2}$.
The detailed study of these examples suggests a clear pattern of how to generalize the procedure and allows to conjecture the general form of the decomposition of the reducible fields that diagonalizes the relevant Lagrangians for an arbitrary value of $n$. These topics are discussed in Section 5.
@@ 2 Reducible ${\cal N}=1$ supermultiplets
Let us briefly summarize the main features of the ${\cal N}=1$ supersymmetric system constructed in [21]. To this end, it is convenient to introduce an auxiliary Fock space, spanned by one set of commuting and one set of anticommuting auxiliary creation and annihilation operators
@@
Let us consider the bosonic fields first, where one has a physical field $|\phi\rangle$ and a set of auxiliary fields. The fields are of the type $|X^{(p,q)}\rangle$, where $p$ is the number of the commuting creation operators, and $q$, which can be either $1$ or 00, is the number of anticommuting creation operators. Obviously, the fields are symmetrical with respect to $p$-indices and no symmetry between $p$-indices and the $q$-index is implied.
@@
The system described above is a particular example of a generalized triplet [13], which contains both totally symmetrical and mixed symmetry reducible representations of of the Poincaré group. The field content (2.2) and the explicit form of the Lagrangian (2) are dictated by the requirement of ${\cal N}=1$ supersymmetry, as we shall see at the end of the present Section.
@@
It can be shown (see [21] for the details), that one can use the gauge invariance (2.6) and the equations of motion in order to eliminate all on-shell fields in (2.2)–(2.3), except $|\phi\rangle$. This vector obeys the massless Klein-Gordon equation, is transverse with respect to both sets of its indices and contains the totally symmetric representations of the Poincaré group of type $Y(n+1,0)$, fields with mixed symmetry of type $Y(n,1)$, as well as their traces. Therefore, one obtains irreducible massless symmetric fields with spins $n+1,n-1,...,1$ or 00 and irreducible massless mixed symmetry fields of types $Y(r,1)$ with $r=n,n-2,...,1$ or 00 (i.e., a vector).
@@
where $|M\rangle=|B\rangle-i|A\rangle$.
The states in the fermionic sector contain only the bosonic oscillators, defined in (2.1). The fields, as well as a Fock vacuum, carry the corresponding $d$dimensional spinorial index, which will not be written explicitly. The field content is
@@
where $|\xi^{(n-1,0)}\rangle$ is a parameter. Again, one can show, that it is possible to gauge away the non-physical fields $|\chi\rangle$ and $|\Sigma\rangle$, and one is left on-shell with the physical field $|\Psi\rangle$. This field obeys the massless Dirac equation, is transverse and describes a massless field with spin $n+\frac{1}{2}$ as well as fields with lower spins $n-\frac{1}{2},...,\frac{1}{2}$, (i.e., irreducible representations of the Poincaré group) since no zero gamma–trace condition is implied.
@@
$\delta\langle A|$ $=$ $\delta\langle E|=\delta\langle F|=0$
which leave the sum of the Lagrangians (2) and (2) invariant. One can check that the supersymmetry algebra close on-shell in the dimensions $d=4,6$ and $10$. Expressing the auxiliary fields via their own equations of motion (2.8), one can can write the supersymmetry transformation rules for fermions as
@@
As we mentioned earlier, the propagating bosonic and fermionic degrees of freedom are contained in $|\phi\rangle$ and $|\Psi\rangle$ respectively, the rest of the fields in (2.2), (2.3) and (2.10) are auxiliary. For $n=0$ the spectrum contains a vector field in the bosonic sector and a spin $\frac{1}{2}$ field in the fermionic sector, i.e., one has an on-shell ${\cal N}=1$ Maxwell supermultiplet in $d=4,6$ and $10$ dimensions. The choice $n=1$ gives ${\cal N}=1$ on-shell supergravity and matter supermultiplets, listed in the Introduction. This pattern (i.e., supermultiplets with $n$ commuting and one anticommuting creation operator in the bosonic field and $n$ commuting creation operators in the fermionic field) holds for an arbitrary value of $n$. The set of auxiliary fields in (2.2), (2.3) and (2.10) is further determined by the requirement of gauge invariance and by ${\cal N}=1$ supersymmetry of the total Lagrangian.
Let us also note that the closure of the supersymmetry algebra, as well as the consistency of the equations (2.8) with the transformation rules (2.13)–(2.14), requires the use of fermionic field equations, derived from the Lagrangian (2). This should have been expected since we are considering on–shell supersymmetry.
@@ 3.1 Flat space-time
For the fermionic fields, the first two steps of the diagonalization procedure are not needed and one starts from the third step. Similarly to the case of the bosonic fields, we introduce a spin $\frac{3}{2}$ field $\tilde{\Psi}_{\mu}$ and a gauge invariant spin $\frac{1}{2}$ field $\tilde{\Psi}$ as
@@
where the coefficient in front of the last term in the definition of the field $\tilde{\Psi}_{\mu}$ is chosen in such a way that the Lagrangian (3.9) written in terms of the new fields is a sum of Fang – Fronsdal Lagrangians for spins $\frac{3}{2}$ and $\frac{1}{2}$
@@
Finally, using the equations (3.6), (3.7), (3.11), as well as (2.14) and (2.15), it is straightforward to rewrite the supersymmetry transformations in terms of irreducible modes. From the field content given above, one can see that this system describes: a linearized ${\cal N}=1$ supergravity and the linear multiplet in case of $d=4$; a linearized ${\cal N}=1$ supergravity and $(1,0)$ tensor multiplet in case of $d=6$; and a linearized ${\cal N}=1$ supegravity in case of $d=10$ dimensions [21].
@@ 3.2 Deformation to $AdS$ space
In order to perform a deformation of the bosonic (2) and fermionic (2) Lagrangians to an $AdS_{d}$ background, one can proceed in two different ways. The first option is to start with the already decomposed system described by a sum of the Lagrangians (3.5), (3.8) and (3.12). Then we perform its deformation to $AdS_{d}$ space by replacing the usual partial derivatives with the covariant ones and adding the corresponding mass-like terms. The Lagrangians for totally symmetric fermionic and bosonic fields can be directly read off from the equations (A) and (A) which we give here for convenience
@@
The mass–like terms for the spin $2$ and spin $\frac{3}{2}$ fields are determined from the requirement that all gauge symmetries of Lagrangians (2) and (2) are kept intact. The mass–like terms for the gauge invariant scalar and spin $\frac{1}{2}$ fields in (3.2) and (3.2) are taken so that they saturate the unitarity bounds for the $AdS_{d}$ isometry group $SO(d-1,2)$. These terms can also be obtained from diagonalization of the spin $2$ and spin $\frac{3}{2}$ triplets on $AdS_{d}$.
@@
where $H_{\mu_{1}\mu_{2}\mu_{3}}=\partial_{[\mu_{1}}{\cal B}_{\mu_{2}\mu_{3}]}$. In the case of $d=4$, one has another option for $AdS$ deformation. Namely, one can first dualize the field ${\cal B}_{\mu\nu}$ to a scalar $a(x)$ as $\partial_{\mu}a\sim\epsilon_{\mu\nu\rho\sigma}H^{\nu\rho\sigma}$ and then perform the $AdS_{4}$ deformation to a conformal scalar, with the mass–like term being the same as in (3.2) for the scalar field $\Phi$. As a result, one obtains either the linear ${\cal N}=1$ supermultiplet [40][41], or the Wess–Zumino supermultiplet, where the spin $\frac{1}{2}$ field being $\tilde{\Psi}$. The requirement of ${\cal N}=1$ supersymmetry for the case of four dimensions555 See [42][43] for a detailed discussion of various properties of the Wess–Zumino supermultiplet on $AdS_{4}$. fixes the mass–like terms for the spin 00 and the spin $\frac{1}{2}$ fields as given in (3.2) and (3.2). Finally, the fields ${\Phi}_{\mu\nu}$ and $\tilde{\Psi}_{\mu}$, in turn form a supermultiplet of the ${\cal N}=1$ supergravity on $AdS_{4}$, constructed in [44].
@@
Using the commutation relations given in the Appendix one can check that Lagrangian (3.2) is invariant under the gauge transformations
@@
$\delta C_{\nu}=-\Box\lambda_{\nu}+\frac{1}{L^{2}}(d-1)\rho_{\nu}+\nabla_{\nu}\tau,$ (3.17)